# Grounded Question-Answering in Long Egocentric Videos

### v3

@@ Appendix B Additional EgoTimeQA Statistics
We offer additional statistical details about our EgoTimeQA. In Fig. 5(a), we present the question distribution based on their first four words. Most questions begin with “what”, inquiring about objects or actions. Others start with “where”, “did”, “how”, etc., showcasing their diversity. In Fig. 5(b) and 5(c), we present the distribution of the top 30 correct and incorrect answers, respectively. Fig. 5(d) presents a histogram of the duration of temporal windows in EgoTimeQA, with the majority falling within the 0-3 second range. Fig. 5(e), 5(f), and 5(g) shows the word count distributions for questions, correct answers, and incorrect answers, respectively. Correct and incorrect answers have similar distributions, averaging 3.2 and 2.8 words respectively, while questions, averaging 6.6 words, indicate their greater complexity.
@@ Appendix E Limitations and Future Work
From the experiments and analysis presented in the main paper and this supplementary material, we can draw the following observations. First, the performance of our method is closely tied to the quality of video features and training data. Enhancements in video features, particularly through improved visual-language alignment, and the inclusion of more training data, could lead to future improvements. Second, despite our efforts in designing appropriate prompts and rigorously filtering data, biases and inaccuracies still exist in the LLM-generated data. Third, our method faces challenges in fine-grained perception tasks, for example, object recognition and counting, particularly in complex environments. The adoption of object-centric features, for example, those for tracking and counting techniques could enhance performance in these areas. Finally, processing long egocentric videos demands significant computational resources. Future research should explore the use of memory networks for compressing video features and developing more efficient models that maintain accuracy.

### v4

@@ Appendix B Additional EgoTimeQA Statistics
We offer additional statistical details about our EgoTimeQA. In Fig. 6(a), we present the question distribution based on their first four words. Most questions begin with “what”, inquiring about objects or actions. Others start with “where”, “did”, “how”, etc., showcasing their diversity. In Fig. 6(b) and 6(c), we present the distribution of the top 30 correct and incorrect answers, respectively. Fig. 6(d) presents a histogram of the duration of temporal windows in EgoTimeQA, with the majority falling within the 0-3 second range. Fig. 6(e), 6(f), and 6(g) shows the word count distributions for questions, correct answers, and incorrect answers, respectively. Correct and incorrect answers have similar distributions, averaging 3.2 and 2.8 words respectively, while questions, averaging 6.6 words, indicate their greater complexity.
@@ Appendix E Limitations and Future Work
From the experiments and analysis presented in the main paper and this supplementary material, we can draw the following observations. First, the performance of our method is closely tied to the quality of video features and training data. Enhancements in video features, particularly through improved visual-language alignment, and the inclusion of more training data, could lead to future improvements. Second, despite our efforts in designing appropriate prompts and rigorously filtering data, biases and inaccuracies still exist in the LLM-generated data. Third, our method faces challenges in fine-grained perception tasks, for example, object recognition and counting, particularly in complex environments. The adoption of object-centric features, for example, those for tracking and counting techniques could enhance performance in these areas. Fourth, processing long egocentric videos demands significant computational resources. Future research should explore the use of memory networks for compressing video features and developing more efficient models that maintain accuracy. Finally, a query may relate to multiple video segments, but Ego4D-NLQ and QA-Ego4D assume only one is relevant per question. We advocate for loosening this assumption, as it is typical for multiple personal experiences to be triggered by a single query. Considering the difficulty of labeling multiple temporal windows for one query, a practical solution is to employ a well-trained Vision-Language Model (VLM) to identify potential candidates for further confirmation via human review. Subsequently, our method can be applied to generate QA pairs for each confirmed candidate.